# Chronos: Learning the Language of Time Series

### v1

@@ 6.3 Data
Our findings underscore that training larger models on a large corpus of time series data yields excellent in-domain and zero-shot performance. Nevertheless, in contrast to NLP, high-quality public time series data remains limited. This poses a dilemma when training models on a large corpus of diverse datasets — selecting more datasets for training leaves fewer for zero-shot evaluation. The time series community would benefit greatly from the availability of larger time series datasets that could be used to develop and improve pretrained model such as Chronos. There have been some recent efforts on building large-scale time series datasets (Emami et al., 2023; Liu et al., 2023) for specific domains, albeit further investment is needed.
@@ Acknowledgements
We are grateful to our fellow researchers who have contributed to this work with insightful discussions and valuable feedback, including but not limited to Laurent Callot, Baris Kurt, Valentin Flunkert, David Salinas, Boran Han, Xiaoyong Jin, Luke Huan, Youngsuk Park, Gaurav Gupta, Karthick Gopalswamy, Tim Januschowski, Jan Gasthaus, Bing Xiang, Kashif Rasul, Mononito Goswami and Gerald Woo.
@@ Appendix A Algorithms
Algorithm 1 and algorithm 2 present the pseudocode for TSMix and KernelSynth, respectively.
Algorithm 1 TSMix: Time Series Mixup
@@ Appendix C Baselines
We consider a total of 14 baseline methods for benchmarking Chronos. Local statistical baselines were AutoETS, AutoARIMA, Naive, Seasonal Naive, and AutoTheta (Assimakopoulos & Nikolopoulos, 2000); for these, we rely on implementations in the StatsForecast library (Garza et al., 2022). For task-specific deep learning architectures, DeepAR (Salinas et al., 2020), PatchTST (Nie et al., 2023), TFT (Lim et al., 2021), DLinear (Zeng et al., 2023), and WaveNet (Oord et al., 2016), we based evaluations on the implementations in GluonTS (Alexandrov et al., 2020). However, N-BEATS (Oreshkin et al., 2020) and N-HiTS (Challu et al., 2023), experiments are based on implementations in the NeuralForecast (Olivares et al., 2022) library. Finally, we use reference implementations of ForecastPFN [^7] https://github.com/abacusai/ForecastPFN (Dooley et al., 2023) and GPT4TS [^8] https://github.com/DAMO-DI-ML/NeurIPS2023-One-Fits-All (One-Fits-All) (Zhou et al., 2023a).
Both training and inference for WaveNet, GPT4TS, and ForecastPFN were conducted on NVIDIA V100 GPUs with 16GB memory on AWS EC2 p3 instances. All other baselines are both trained on the CPU on Intel-based EC2 instances. Task-specific deep learning baselines not based on large language models (DeepAR, PatchTST, TFT, DLinear, WaveNet, N-BEATS, and N-HiTS) are trained and evaluated three times and their performance averaged in order to account for high variance inherent in their optimization.
Statistical baselines (AutoETS, AutoARIMA, AutoTheta and SeasonalNaive) are used with their default hyperparameters in StatsForecast, but with season lengths implied by their frequencies. For example, daily frequency data had season length set to 7, hourly data 24, etc. For this heuristic, we use the helper function get_seasonality from GluonTS.
In PatchTST and DLinear, we experiment with two loss functions: original losses aimed at point forecasting (L1 or L2 loss) as well as default probabilistic forecasting heads used in their GluonTS implementations, where loss is set to the negative Student’s $t$ log likelihood of the forecast horizon. Due to the consistently superior performance, our final results include probabilistic versions of PatchTST and DLinear only.
Default hyperparameter configurations provided in baseline implementations are kept as is, and no dataset specific or global hyperparameter tuning is performed. GluonTS-based implementations sre optimized with a batch size of 128, for a time limit of 4 hours and early stopping patience of 200 epochs. For GPT4TS, we set the context length equal to a multiple of the prediction length, with the multiplier depending on the frequency of the dataset (Table 3). We use the MASE loss function for fine-tuning in GPT4TS due to its superior performance. A summary of the baseline models used along with details of hyperparameter values is provided in Table 4.
@@
SeasonalNaive Local StatsForecast Yes
N/A
AutoETS Local StatsForecast Yes
$C=2500$
AutoARIMA Local StatsForecast Yes
$C=1000$
AutoTheta Local StatsForecast Yes
$C=2500$
DeepAR Task-specific GluonTS Yes
$d_{h}=40,n_{L}=2$
TFT Task-specific GluonTS Yes
$d_{h}=32,n_{H}=4$
PatchTST Task-specific GluonTS Yes
Patch length: 16, Stride: 8, $d_{h}=32,n_{L}=2,n_{H}=4$
DLinear Task-specific GluonTS Yes
Kernel size: 25, $d_{h}=20$
WaveNet Task-specific GluonTS Yes
Residual channels: 24, Skip channels: 3
N-BEATS Task-specific NeuralForecast No
Input size multiplier: 5
N-HiTS Task-specific NeuralForecast No
Input size multiplier: 5
GPT4TS Task-specific Reference No
Fine-tuning epochs: 100, cos: 1, tmax: 10, $n_{L}=6,\eta=10^{-3}$, with pretrained GPT-2 weights
ForecastPFN Zero-shot Reference No
$C=100$ (as in the released pretrained model)
@@ Appendix E Additional Results
Pretrained Models (In Domain) Task Specific Models Local Models
[…]
@@
Electricity (15 Min.) 0.076 0.076 0.081 0.081 0.077 0.082 0.090 0.091 0.189 0.079 0.081 0.084 - 0.229 - 0.117 0.279
Electricity (Hourly) 0.102 0.115 0.107 0.092 0.121 0.089 0.106 0.109 0.125 0.095 0.128 0.127 0.129 0.198 0.126 0.147 0.363
Electricity (Weekly) 0.064 0.067 0.077 0.078 0.069 0.069 0.116 0.105 0.106 0.146 0.098 0.097 0.151 0.146 0.138 0.198 0.198
KDD Cup 2018 0.273 0.272 0.294 0.271 0.359 0.252 0.330 0.280 0.571 0.312 0.302 0.315 2.266 0.521 0.528 0.556 -
London Smart Meters 0.423 0.426 0.430 0.433 0.431 0.346 0.405 0.374 0.365 0.369 0.358 0.357 - 0.660 - 0.541 0.731
M4 (Daily) 0.021 0.021 0.021 0.021 0.020 0.023 0.023 0.023 0.023 0.024 0.022 0.022 0.027 0.024 0.023 0.028 0.028
M4 (Hourly) 0.025 0.028 0.026 0.025 0.038 0.027 0.038 0.046 0.033 0.038 0.040 0.045 0.066 0.041 - 0.048 0.166
M4 (Monthly) 0.102 0.103 0.104 0.104 0.110 0.095 0.101 0.107 0.097 0.111 0.094 0.093 0.100 0.098 - 0.146 0.140
M4 (Weekly) 0.038 0.039 0.041 0.042 0.040 0.039 0.046 0.045 0.051 0.044 0.039 0.040 0.052 0.053 0.050 0.063 0.063
Pedestrian Counts 0.198 0.203 0.234 0.241 0.175 0.257 0.229 0.248 0.261 0.247 0.254 0.241 0.619 1.818 0.340 0.319 0.814
Rideshare 0.141 0.140 0.141 0.135 0.141 0.135 0.130 0.184 0.134 0.159 0.152 0.172 0.154 0.138 0.157 0.186 -
Taxi (30 Min.) 0.267 0.273 0.312 0.311 0.335 0.363 0.395 0.347 0.382 0.335 0.306 0.305 - 0.456 - 0.471 0.741
Temperature-Rain 0.663 0.670 0.686 0.704 0.669 0.804 0.718 0.708 0.670 0.848 0.780 0.798 1.182 1.060 0.869 1.424 -
Uber TLC (Daily) 0.096 0.097 0.099 0.106 0.099 0.100 0.110 0.126 0.111 0.106 0.116 0.108 0.167 0.190 0.151 0.231 0.231
Uber TLC (Hourly) 0.155 0.154 0.156 0.161 0.161 0.167 0.176 0.168 0.179 0.234 0.166 0.161 0.462 0.433 0.311 0.299 0.625
[…]
@@
Pretrained Models (In Domain) Task Specific Models Local Models
[…]
@@
Electricity (15 Min.) 0.403 0.409 0.430 0.453 0.406 0.450 0.515 0.637 1.108 0.452 0.579 0.567 0.508 - 0.583 - 0.498 1.270
Electricity (Hourly) 1.457 1.602 1.500 1.370 1.653 1.349 1.528 1.537 1.789 1.369 1.880 1.848 1.487 1.774 2.151 1.715 1.840 4.159
Electricity (Weekly) 1.829 1.888 2.007 2.044 1.879 1.631 2.517 1.929 2.800 2.613 1.975 2.035 1.880 3.086 3.078 3.009 3.037 3.037
KDD Cup 2018 0.663 0.673 0.688 0.656 0.771 0.616 0.779 0.671 1.022 0.695 0.674 0.731 0.737 1.014 1.138 1.023 0.994 -
London Smart Meters 0.843 0.852 0.859 0.868 0.856 0.733 0.832 0.824 0.788 0.799 0.777 0.781 0.794 - 0.966 - 0.966 1.297
M4 (Daily) 3.099 3.109 3.112 3.108 3.058 3.450 3.305 3.306 3.292 3.461 3.143 3.155 5.109 3.270 3.335 3.257 3.278 3.278
M4 (Hourly) 0.927 0.894 0.881 0.879 0.977 0.967 1.215 1.613 1.833 1.867 3.231 3.457 1.511 1.604 2.458 - 1.193 11.608
M4 (Monthly) 0.976 0.985 1.000 1.006 1.051 0.962 1.040 1.101 1.009 1.022 0.994 0.942 0.979 0.970 0.966 - 1.260 1.205
M4 (Weekly) 2.201 2.221 2.270 2.309 2.377 1.996 2.346 2.523 2.745 2.429 2.094 1.976 3.040 2.548 2.657 2.373 2.777 2.777
Pedestrian Counts 0.289 0.286 0.304 0.309 0.277 0.339 0.311 0.334 0.364 0.327 0.324 0.315 0.393 0.487 1.275 0.383 0.369 0.842
Rideshare 0.892 0.892 0.880 0.857 0.897 0.827 0.996 0.983 1.067 1.448 0.933 0.919 1.088 0.910 0.970 1.028 1.250 -
Taxi (30 Min.) 0.828 0.849 0.938 0.939 1.032 1.077 1.158 1.070 1.113 1.018 0.950 0.934 1.113 - 1.193 - 1.160 1.768
Temperature-Rain 0.982 0.991 1.013 1.033 0.984 1.250 1.015 1.076 0.994 1.370 1.232 1.343 1.226 1.968 1.945 1.524 2.243 -
Uber TLC (Daily) 0.819 0.836 0.871 0.906 0.846 0.813 0.905 0.938 0.916 0.855 0.877 0.879 0.838 1.228 1.312 1.114 1.378 1.378
Uber TLC (Hourly) 0.727 0.729 0.727 0.743 0.740 0.696 0.703 0.776 0.746 0.778 0.716 0.751 0.754 1.009 1.036 0.982 0.931 1.390
[…]
@@
Pretrained Models (Zero Shot) Task Specific Models Local Models
[…]

### v2

@@ 6.3 Data
Our findings underscore that training larger models on a large corpus of time series data yields excellent in-domain and zero-shot performance. Nevertheless, in contrast to NLP, high-quality public time series data remains limited. This poses a dilemma when training models on a large corpus of diverse datasets — selecting more datasets for training leaves fewer for zero-shot evaluation. The time series community would benefit greatly from the availability of larger time series datasets that could be used to develop and improve pretrained model such as Chronos. There have been some recent efforts on building large-scale time series datasets for specific domains (Emami et al., 2023; Liu et al., 2023) and cross-domain (Borchert et al., 2022), albeit further investment is needed.
@@ Acknowledgements
We are indebted to Stefano Soatto for challenging us to think about the fundamental question regarding language models and time series modeling, ultimately leading to the creation of the present work. We are grateful to our fellow researchers who have contributed to this work with insightful discussions and valuable feedback, including but not limited to Devamanyu Hazarika, Imry Kissos, Laurent Callot, Baris Kurt, Valentin Flunkert, David Salinas, Boran Han, Xiaoyong Jin, Luke Huan, Youngsuk Park, Gaurav Gupta, Karthick Gopalswamy, Tim Januschowski, Jan Gasthaus, Bing Xiang, Kashif Rasul, Juba Nait Saada, Matthias Karlbauer, Mononito Goswami and Gerald Woo.
@@ Appendix A Algorithms
Algorithm 1 and algorithm 2 present the pseudocode for TSMixup and KernelSynth, respectively.
Algorithm 1 TSMixup: Time Series Mixup
@@ Appendix C Baselines
We considered a total of 17 baseline methods for benchmarking Chronos. Local statistical baselines were AutoETS, AutoARIMA, Naive, Seasonal Naive, and AutoTheta (Assimakopoulos & Nikolopoulos, 2000); for these, we relied on implementations in the StatsForecast library (Garza et al., 2022). For task-specific deep learning architectures, DeepAR (Salinas et al., 2020), PatchTST (Nie et al., 2023), TFT (Lim et al., 2021), DLinear (Zeng et al., 2023), and WaveNet (Oord et al., 2016), we based evaluations on the implementations in GluonTS (Alexandrov et al., 2020). However, N-BEATS (Oreshkin et al., 2020) and N-HiTS (Challu et al., 2023), experiments were based on implementations in the NeuralForecast (Olivares et al., 2022) library. Finally, we used reference implementations of ForecastPFN [^7] https://github.com/abacusai/ForecastPFN (Dooley et al., 2023), GPT4TS [^8] https://github.com/DAMO-DI-ML/NeurIPS2023-One-Fits-All (One-Fits-All) (Zhou et al., 2023a), LLMTime [^9] https://github.com/ngruver/llmtime (Gruver et al., 2023), Lag-Llama [^10] https://github.com/time-series-foundation-models/lag-llama (Rasul et al., 2023), and Moirai-1.0-R [^11] https://github.com/SalesforceAIResearch/uni2ts (Woo et al., 2024).
WaveNet and GPT4TS models were trained on AWS EC2 p3.2xlarge instances which have 1 NVIDIA V100 GPUs with 16GB VRAM. All other baselines were trained on the CPU on Intel-based EC2 instances. Task-specific deep learning baselines not based on large language models (DeepAR, PatchTST, TFT, DLinear, WaveNet, N-BEATS, and N-HiTS) were trained and evaluated three times and their performance averaged in order to account for high variance inherent in their optimization.
For inference, we used EC2 CPU instances for local models, N-HiTS, and N-BEATS. The p3.2xlarge instance (1 $\times$ V100 16GB) was used for inference for other task-specific deep learning models and pretrained models such as Lag-Llama, Moirai-1.0-R, and ForecastPFN. Since LLMTime uses a Llama-2 70B model which has significantly larger compute requirements, LLMTime inference was performed on the p3dn.24xlarge AWS EC2 instance with 8 NVIDIA V100 32GB GPUs.
@@
Statistical baselines (AutoETS, AutoARIMA, AutoTheta and SeasonalNaive) were used with their default hyperparameters in StatsForecast, but with season lengths implied by their frequencies. For example, daily frequency data had season length set to 7, hourly data 24, and so on. For this heuristic, we used the helper function get_seasonality from GluonTS.
Unless otherwise specified, the default hyperparameter configurations provided in baseline implementations were kept as is, and no dataset specific or global hyperparameter tuning was performed. GluonTS-based implementations were optimized with a batch size of 128, for a time limit of 4 hours and early stopping patience of 200 epochs. In PatchTST and DLinear, we experimented with two loss functions: original losses aimed at point forecasting (L1 or L2 loss) as well as default probabilistic forecasting heads used in their GluonTS implementations, where the loss is set to the negative Student’s-$t$ log likelihood of the forecast horizon. Due to the consistently superior performance, our final results include the probabilistic versions of PatchTST and DLinear only. For GPT4TS, we set the context length equal to a multiple of the prediction length, with the multiplier depending on the frequency of the dataset (Table 3). We used the MASE loss function for fine-tuning in GPT4TS due to its superior performance.
For LLMTime, we experimented only with the Llama-2 70B due to the prohibitively high costs of running the benchmark through OpenAI APIs. We used the same hyperparameters as used in the Monash experiment in the original paper (Gruver et al., 2023) with a few notable differences. We set the context length to 512, same as for Chronos models, instead of 500. During our experiments, we observed that the default hyperparameters may lead to a significant drop in the scale of the last prediction on some datasets. To alleviate this issue, we set the STEP_MULTIPLIER to 1.4 (instead of 1.2) and increased the prediction length by 1 (this extra prediction is removed before computing the metrics). The inference time for LLMTime (Llama-2 70B) is $\approx$0.8 seconds per observation on p3dn.24xlarge. As an example, this will take 92 hours to generate all the predictions on the Traffic dataset (862 time series, 24 as prediction length, 20 samples). Due to the very high compute cost, we skip the evaluation of LLMTime on some large datasets.
A summary of the baseline models used along with details of hyperparameter values is provided in Table 4.
@@
Naive Local StatsForecast Yes N/A
SeasonalNaive Local StatsForecast Yes N/A
AutoETS Local StatsForecast Yes $C=2500$
AutoARIMA Local StatsForecast Yes $C=1000$
AutoTheta Local StatsForecast Yes $C=2500$
DeepAR Task-specific GluonTS Yes $d_{h}=40,n_{L}=2$
TFT Task-specific GluonTS Yes $d_{h}=32,n_{H}=4$
PatchTST Task-specific GluonTS Yes Patch length: 16, Stride: 8, $d_{h}=32,n_{L}=2,n_{H}=4$
DLinear Task-specific GluonTS Yes Kernel size: 25, $d_{h}=20$
WaveNet Task-specific GluonTS Yes Residual channels: 24, Skip channels: 3
N-BEATS Task-specific NeuralForecast No Input size multiplier: 5
N-HiTS Task-specific NeuralForecast No Input size multiplier: 5
GPT4TS Task-specific Reference No Fine-tuning epochs: 100, cos: 1, tmax: 10, $n_{L}=6,\eta=10^{-3}$, with pretrained GPT-2 weights
ForecastPFN Pretrained Reference No $C=100$ (as in the released pretrained model)
[…]
Lag-Llama Pretrained Reference Yes $C=32$
Moirai-1.0-R Pretrained Reference Yes $C=1024$, Patch length: selected by dataset-specific validation
@@ Appendix E Additional Results
Pretrained Models (In Domain) Pretrained Models (Other) Task Specific Models Local Models
@@
Electricity (15 Min.) 0.076 0.076 0.081 0.081 0.077 0.319 0.105 0.103 0.082 0.090 0.091 0.189 0.079 0.081 0.084 - 0.229 - 0.117 0.279
Electricity (Hourly) 0.102 0.115 0.107 0.092 0.121 0.104 0.122 0.116 0.089 0.106 0.109 0.125 0.095 0.128 0.127 0.129 0.198 0.126 0.147 0.363
Electricity (Weekly) 0.064 0.067 0.077 0.078 0.069 0.147 0.110 0.159 0.069 0.116 0.105 0.106 0.146 0.098 0.097 0.151 0.146 0.138 0.198 0.198
KDD Cup 2018 0.273 0.272 0.294 0.271 0.359 0.369 0.286 0.277 0.252 0.330 0.280 0.571 0.312 0.302 0.315 2.266 0.521 0.528 0.556 -
London Smart Meters 0.423 0.426 0.430 0.433 0.431 0.384 0.358 0.350 0.346 0.405 0.374 0.365 0.369 0.358 0.357 - 0.660 - 0.541 0.731
M4 (Daily) 0.021 0.021 0.021 0.021 0.020 0.043 0.023 0.023 0.023 0.023 0.023 0.023 0.024 0.022 0.022 0.027 0.024 0.023 0.028 0.028
M4 (Hourly) 0.025 0.028 0.026 0.025 0.038 0.111 0.025 0.022 0.027 0.038 0.046 0.033 0.038 0.040 0.045 0.066 0.041 - 0.048 0.166
M4 (Monthly) 0.102 0.103 0.104 0.104 0.110 0.153 0.102 0.100 0.095 0.101 0.107 0.097 0.111 0.094 0.093 0.100 0.098 - 0.146 0.140
M4 (Weekly) 0.038 0.039 0.041 0.042 0.040 0.078 0.049 0.046 0.039 0.046 0.045 0.051 0.044 0.039 0.040 0.052 0.053 0.050 0.063 0.063
Pedestrian Counts 0.198 0.203 0.234 0.241 0.175 0.262 0.274 0.259 0.257 0.229 0.248 0.261 0.247 0.254 0.241 0.619 1.818 0.340 0.319 0.814
Rideshare 0.141 0.140 0.141 0.135 0.141 0.158 0.164 0.159 0.135 0.130 0.184 0.134 0.159 0.152 0.172 0.154 0.138 0.157 0.186 -
Taxi (30 Min.) 0.267 0.273 0.312 0.311 0.335 0.357 0.513 0.368 0.363 0.395 0.347 0.382 0.335 0.306 0.305 - 0.456 - 0.471 0.741
Temperature-Rain 0.663 0.670 0.686 0.704 0.669 0.717 0.655 0.685 0.804 0.718 0.708 0.670 0.848 0.780 0.798 1.182 1.060 0.869 1.424 -
Uber TLC (Daily) 0.096 0.097 0.099 0.106 0.099 0.176 0.116 0.108 0.100 0.110 0.126 0.111 0.106 0.116 0.108 0.167 0.190 0.151 0.231 0.231
Uber TLC (Hourly) 0.155 0.154 0.156 0.161 0.161 0.176 0.176 0.167 0.167 0.176 0.168 0.179 0.234 0.166 0.161 0.462 0.433 0.311 0.299 0.625
Agg. Relative Score 0.574 0.589 0.610 0.607 0.624 0.937 0.688 0.667 0.601 0.676 0.689 0.734 0.697 0.656 0.664 1.076 1.083 0.876 1.000 1.433
Avg. Rank 3.400 4.867 6.600 6.267 6.933 14.267 10.800 9.200 6.133 9.600 10.600 10.267 10.400 8.000 8.333 16.700 14.933 16.367 17.400 18.933
@@
Pretrained Models (In Domain) Pretrained Models (Other) Task Specific Models Local Models
@@
Electricity (15 Min.) 0.403 0.409 0.430 0.453 0.406 1.169 0.706 0.626 0.450 0.515 0.637 1.108 0.452 0.579 0.567 0.508 - 0.583 - 0.498 1.270
Electricity (Hourly) 1.457 1.602 1.500 1.370 1.653 1.573 1.713 1.665 1.349 1.528 1.537 1.789 1.369 1.880 1.848 1.487 1.774 2.151 1.715 1.840 4.159
Electricity (Weekly) 1.829 1.888 2.007 2.044 1.879 2.979 2.848 2.759 1.631 2.517 1.929 2.800 2.613 1.975 2.035 1.880 3.086 3.078 3.009 3.037 3.037
KDD Cup 2018 0.663 0.673 0.688 0.656 0.771 0.844 0.661 0.656 0.616 0.779 0.671 1.022 0.695 0.674 0.731 0.737 1.014 1.138 1.023 0.994 -
London Smart Meters 0.843 0.852 0.859 0.868 0.856 0.792 0.770 0.755 0.733 0.832 0.824 0.788 0.799 0.777 0.781 0.794 - 0.966 - 0.966 1.297
M4 (Daily) 3.099 3.109 3.112 3.108 3.058 8.038 3.443 3.376 3.450 3.305 3.306 3.292 3.461 3.143 3.155 5.109 3.270 3.335 3.257 3.278 3.278
M4 (Hourly) 0.927 0.894 0.881 0.879 0.977 3.807 1.209 0.951 0.967 1.215 1.613 1.833 1.867 3.231 3.457 1.511 1.604 2.458 - 1.193 11.608
M4 (Monthly) 0.976 0.985 1.000 1.006 1.051 2.090 1.033 1.005 0.962 1.040 1.101 1.009 1.022 0.994 0.942 0.979 0.970 0.966 - 1.260 1.205
M4 (Weekly) 2.201 2.221 2.270 2.309 2.377 5.658 2.466 2.391 1.996 2.346 2.523 2.745 2.429 2.094 1.976 3.040 2.548 2.657 2.373 2.777 2.777
Pedestrian Counts 0.289 0.286 0.304 0.309 0.277 0.342 0.355 0.330 0.339 0.311 0.334 0.364 0.327 0.324 0.315 0.393 0.487 1.275 0.383 0.369 0.842
Rideshare 0.892 0.892 0.880 0.857 0.897 0.891 0.911 0.899 0.827 0.996 0.983 1.067 1.448 0.933 0.919 1.088 0.910 0.970 1.028 1.250 -
Taxi (30 Min.) 0.828 0.849 0.938 0.939 1.032 1.069 1.374 1.089 1.077 1.158 1.070 1.113 1.018 0.950 0.934 1.113 - 1.193 - 1.160 1.768
Temperature-Rain 0.982 0.991 1.013 1.033 0.984 1.031 0.963 0.988 1.250 1.015 1.076 0.994 1.370 1.232 1.343 1.226 1.968 1.945 1.524 2.243 -
Uber TLC (Daily) 0.819 0.836 0.871 0.906 0.846 1.289 0.935 0.878 0.813 0.905 0.938 0.916 0.855 0.877 0.879 0.838 1.228 1.312 1.114 1.378 1.378
Uber TLC (Hourly) 0.727 0.729 0.727 0.743 0.740 0.711 0.727 0.717 0.696 0.703 0.776 0.746 0.778 0.716 0.751 0.754 1.009 1.036 0.982 0.931 1.390
Agg. Relative Score 0.726 0.736 0.751 0.752 0.763 1.141 0.855 0.806 0.740 0.821 0.842 0.939 0.864 0.854 0.861 0.871 0.983 1.129 0.941 1.000 1.484
Avg. Rank 4.000 5.533 6.333 7.000 7.400 13.467 11.267 8.933 5.133 10.400 11.933 13.533 11.467 9.000 9.333 11.600 15.833 16.533 16.967 16.133 19.200
@@
Pretrained Models (Zero Shot) Pretrained Models (Other) Task Specific Models Local Models
@@
LLMTime
Lag-Llama
Moirai-1.0-R (Base)
Moirai-1.0-R (Large)